# Non-Gaussian enhancement of precision in quantum batteries

### v1

@@ Appendix F Effects of truncation
Appendix G Multimode case
When the battery system comprises multiple modes, the possibility to perform entangling operations between them is unlocked. This naturally raises the question of whether such operations can offer an advantage in the relative precision of the stored energy. In the most general case, each of the modes may have different oscillating frequencies, and a comment in this regard is made in Friis and Huber (2018) on the fact that Gaussian passive optics are sufficient to mix the modes in such a way that photons are stored in the lower frequency modes, implying a gain in energy precision.
However, in our scenario the response of the harvester is uniquely dependent on the photon number statistics (thus equivalent to the case where the battery modes all oscillate at the same frequency $\omega$). In that case, it can be shown that through Gaussian global (entangling) unitaries only, our figure of merit $\Gamma$ remains constant. On the contrary, non-Gaussian states feature signal-to-noise ratios that are tunable by passive, entangling operations, allowing for some extra advantage with respect to the Gaussian case when the correct beam-splitting angle is applied. In this sense, our figure-of-merit constitutes a signature of non-Gaussianity, because the dependency of $\Gamma$ on the parameters of an entangling operation, such as a beam splitter, is an exclusive property of non-Gaussian states.
@@ Appendix G Multimode case
In Fig.7 we show the value of $\Gamma$ for the optimal two-mode one-photon added state, both in the separable case (yielding the worst ratio) and the maximally entangled one. The shaded area represents the gain given by the application of a $\theta=\pi/4$ beam splitter. Note that, for low temperatures, entanglement plays no role since, as mentioned above, close to the pure state regime, Fock states are always optimal in terms of photon number uncertainty. Finally, we address the question of whether this advantage is extensive with the number of modes. In Fig.8 we show the optimal signal-to-noise ratio $\Gamma$ (upon tuning of the squeezing, beam-splitting, phase-shifting and displacement parameters) attainable for multimode batteries of increasing degrees of non-Gaussianity. From the results, we conclude that the hierarchy in extractable signal-to-noise ratios established by the degree of non-Gaussianity in the single mode case is a property that can be generalized to larger $N$ , and the existent increasing trend of $\Gamma$ with $N$ suggests that bigger and more complex non-gaussian battery systems hold a promising potential as highly precise energy storage devices.

### v2

@@ Appendix G Appendix G: Multimode case
Finally, we address the question of whether this advantage is extensive with the number of modes. In Fig.8 we show the optimal signal-to-noise ratio $\Gamma$ (upon tuning of the squeezing, beam-splitting, phase-shifting and displacement parameters) attainable for multimode batteries of increasing degrees of non-Gaussianity. From the results, we conclude that the hierarchy in extractable signal-to-noise ratios established by the degree of non-Gaussianity in the single mode case is a property that can be generalized to larger $N$, and the existent increasing trend of $\Gamma$ with $N$ suggests that bigger and more complex non-gaussian battery systems hold a promising potential as highly precise energy storage devices.
@@
Appendix H Appendix H
H.1 1. SNR under lossy/noisy detection
Let $\rho$ be the photonic output state with $\delta N=\langle N\rangle-N_{0}$ and $(\Delta N)^{2}=\mathrm{Var}_{\rho}(N)$, and define the state-level SNR $\Gamma=\delta N/\Delta N$. We model the detection/readout as a linear loss channel of efficiency $\eta\in[0,1]$ (binomial thinning of photon counts) followed by additive, classical readout noise $B$ that is independent of $N$, with mean $\beta$ and variance $\sigma_{B}^{2}$. Applying the same map to both the initial and the charged state gives
$\langle N_{\rm det}\rangle$ $=\eta\,\langle N\rangle,$ (65)
$\mathrm{Var}(N_{\rm det})$ $=\eta^{2}\,\mathrm{Var}(N)\;+\;\eta(1-\eta)\,\langle N\rangle\;+\;\sigma_{B}^{2},$ (66)
so that the measured SNR is
$\Gamma_{\rm det}=\frac{\delta N_{\rm det}}{\sqrt{\mathrm{Var}(N_{\rm det})}}=\frac{\eta\,\delta N}{\sqrt{\eta^{2}\Delta N^{2}+\eta(1-\eta)\langle N\rangle+\sigma_{B}^{2}}}.$ (67)
Since $\eta(1-\eta)\langle N\rangle+\sigma_{B}^{2}\geq 0$, one has $\mathrm{Var}(N_{\rm det})\geq\eta^{2}\Delta N^{2}$, which implies the monotonicity (data-processing) inequality
$\Gamma_{\rm det}\ \leq\ \Gamma,$ (68)
[…]
For $T=1\,\mathrm{K}$ ($\nu\simeq 1.313$, $N_{0}\simeq 0.157$), energy budget $\epsilon=5$, overall efficiency $\eta=0.60$, and electronic noise $\sigma_{B}=0.30$, the propagated Gaussian bound gives $\Gamma_{G,\mathrm{det}}\approx 1.84$. An optimised single photon–added state (with $\Gamma\simeq 1.6\,\Gamma_{G}\approx 4.16$) attains $\Gamma_{\mathrm{det}}\approx 2.21>\Gamma_{G,\mathrm{det}}$, thus violating the bound at the electrical readout.
H.2 2. Finite-time TUR for the electrical readout
We evaluate precision on the integrated charge $Q_{\tau}=\int_{0}^{\tau}I(t)\,dt$ delivered to a resistive load at (effective) temperature $T_{e}$ over a window $\tau$ (accumulate-and-sample). The measured SNR is
$\Gamma_{\tau}\ :=\ \frac{\langle Q_{\tau}\rangle-\langle Q_{\tau}\rangle_{0}}{\sqrt{\mathrm{Var}(Q_{\tau})}}\,.$ (69)
In stochastic thermodynamics, the finite-time thermodynamic uncertainty relation (TUR) for an integrated current yields
$\frac{\mathrm{Var}(Q_{\tau})}{\langle Q_{\tau}\rangle^{2}}\ \geq\ \frac{2}{\Sigma_{\tau}}\,,\qquad\Sigma_{\tau}=\frac{1}{k_{B}T_{e}}\int_{0}^{\tau}\langle P_{\rm diss}(t)\rangle\,dt,\quad P_{\rm diss}(t)=I(t)^{2}R,$ (70)
where $\Sigma_{\tau}$ is the total (dimensionless) entropy production in the load during $\tau$ [BaratoSeifert2015_TUR, Gingrich2016_DissipationBounds, DechantSasa2018_EntropicBound, KoyukSeifert2019_TimeDependentTUR, HorowitzGingrich2020_NatPhys_TURReview]. For stationary operation with mean current $\langle I\rangle$, one has $\Sigma_{\tau}=(\langle I\rangle^{2}R/k_{B}T_{e})\,\tau$, so
$\Gamma_{\tau}\ \leq\ \sqrt{\Sigma_{\tau}/2}\ =\ \sqrt{\frac{\langle I\rangle^{2}R\,\tau}{2k_{B}T_{e}}}\,.$ (71)
This ceiling is detector-level and independent of how the optical state is prepared. Optical non-classicality/non-Gaussianity can reduce $\mathrm{Var}(Q_{\tau})$ at fixed mean, thereby pushing $\Gamma_{\tau}$ closer to the same ceiling; our witnesses (Theorems AB) remain state-level and are tested by comparing $\Gamma_{\rm det}$ to detection-propagated bounds (See Appendix H.1 above). The TUR statement only supplies an absolute, dissipation-limited benchmark at the harvester.